# Thermoelectric coefficients of two-flavor quark matter from the Kubo formalism

### v1

@@ V Summary
In this work, we studied the thermoelectric power and the Thomson coefficient of quark matter within the two-flavor NJL model using the Kubo formalism. We have derived the relevant Kubo formulas for thermoelectric coefficients starting from the quantum Liouville equation for the density matrix, assuming the simultaneous presence of a weak external electric field and thermodynamic gradients. The relevant correlation functions are computed for a relativistic quark plasma with full quark propagator and within the $1/N_{c}$ approximation to the multi-loop contributions, which implies that the leading-order contributions to the correlation functions arise from single-loop diagrams. The full Lorentz structure of quark spectral functions is taken into account. Here we used quark spectral functions derived from one meson exchange processes above the Mott transition temperatures LKW15 ; Harutyunyan2017 . Numerical results for the thermopower and the Thomson coefficient were obtained over a range of temperatures and chemical potentials relevant primarily to heavy-ion collisions. The results show that both coefficients increase nearly linearly with temperature, indicating that higher temperatures enhance the system’s ability to generate electric fields in response to a thermal gradient. We also find that these coefficients become divergent in the limit where the chemical potential vanishes: this feature is a direct consequence of the fact that the particle rest frame relative to which heat transfer is defined, becomes ill-defined in this limit. We also estimated the values of electric fields which are generated in quark matter due to the thermoelectric effect (Seebeck effect), and found that the effect is more pronounced at higher temperatures and relatively lower chemical potentials.
@@ Appendix A Solving the Liouville equation
where $\Omega_{l}(t)=-\ln{\rm Tr}{e^{-A_{H}(t)}}$. In Eq. (104) we omit the index $H$ in the operators $T^{00}$ and $N^{0}$ for simplicity (the explicit time-dependence already assumes the Heisenberg representation). Note that, as the expression (A) contains $\dot{a}_{H}(t)$ only in the commutator with $\rho_{0}$, it is sufficient to keep only the term $\dot{A}_{H}(t)$ in $\dot{a}_{H}(t)$, as $\dot{\Omega}_{l}(t)$ is a number, and, therefore, commutes with any operator. We next compute the time derivative of the operator $A_{H}(t)$
@@
therefore, taking $\nu=0$, we obtain $\partial_{0}T^{00}=-\partial_{i}T^{i0}-J_{i}E^{i}$, $\partial_{0}N^{0}=-\partial_{i}N^{i}$, where $F^{0i}=-E^{i}$ is the electric field for fluids at rest. Therefore (we omit the arguments $(\bm{r},t)$ for simplicity)
@@
with $h=\mu+sT/n$ being the enthalpy per particle. As this relation holds for any quasi-equilibrium variations of $p$, $\beta$ and $\alpha$, it can also be applied for the partial derivatives of these parameters. Since we consider a fluid at rest, the gradient of pressure vanishes due to the condition of mechanical equilibrium $\partial_{i}p=0$. Employing this, we arrive at $\partial_{i}\alpha=h\partial_{i}\beta$, and substituting this back into Eq. (109) we finally obtain
@@
Note that $U_{H}(t^{\prime})$ contains the operators of the energy density $T^{00}$ and the net particle density $N^{0}$, which do not contribute to the averages of the current operators according to Curie’s theorem. The latter states that in an isotropic medium the correlations between operators of different rank vanish 1963PhT….16e..70D ; 1963AmJPh..31..558D . Therefore, in the solution (A), the term $U(t^{\prime})$ can be dropped. Within the linear response approximation, we can further ignore the space-time dependence of $\beta(\bm{r}^{\prime},t^{\prime\prime})$ in the integrand of Eq. (A) and cancel it with the term $1/\beta$.

### v2

@@ V Summary
In this work, we studied the thermoelectric power and the Thomson coefficient of quark matter within the two-flavor NJL model using the Kubo formalism. We have derived the relevant Kubo formulas for thermoelectric coefficients starting from the quantum Liouville equation for the density matrix, assuming the simultaneous presence of a weak external electric field and thermodynamic gradients. The relevant correlation functions are computed for a relativistic quark plasma with full quark propagator and within the $1/N_{c}$ approximation to the multiloop contributions, which implies that the leading-order contributions to the correlation functions arise from single-loop diagrams. The full Lorentz structure of quark spectral functions is taken into account. Here we used quark spectral functions derived from one-meson-exchange processes above the Mott transition temperatures Lang et al. (2015); Harutyunyan et al. (2017). Numerical results for the thermopower and the Thomson coefficient were obtained over a range of temperatures and chemical potentials relevant primarily to heavy-ion collisions. The results show that both coefficients increase nearly linearly with temperature, indicating that higher temperatures enhance the system’s ability to generate electric fields in response to a thermal gradient. We also find that these coefficients diverge in the limit of vanishing chemical potential. This behavior is a direct consequence of the fact that the particle rest frame, relative to which heat transport is defined, becomes ill defined in this limit.
We further estimated the magnitude of the electric fields generated in quark matter by the thermoelectric (Seebeck) effect and find that the effect is more pronounced at higher temperatures and relatively low chemical potentials. Finally, we compared our results with previous calculations performed within the RTA. We find that the Seebeck coefficient obtained here is larger than perturbative QCD estimates, while remaining consistent in order of magnitude with earlier NJL-model studies that employed again in the RTA.
@@ Appendix A Solving the Liouville equation
where $\Omega_{l}(t)=-\ln{\rm Tr}{e^{-A_{H}(t)}}$. In Eq. (105) we omit the index $H$ in the operators $T^{00}$ and $N^{0}$ for simplicity (the explicit time dependence already assumes the Heisenberg representation). Note that, as the expression (A) contains $\dot{a}_{H}(t)$ only in the commutator with $\rho_{0}$, it is sufficient to keep only the term $\dot{A}_{H}(t)$ in $\dot{a}_{H}(t)$, as $\dot{\Omega}_{l}(t)$ is a number, and, therefore, commutes with any operator. We next compute the time derivative of the operator $A_{H}(t)$,
@@
therefore, taking $\nu=0$, we obtain $\partial_{0}T^{00}=-\partial_{i}T^{i0}-J_{i}E^{i}$, $\partial_{0}N^{0}=-\partial_{i}N^{i}$, where $F^{0i}=-E^{i}$ is the electric field for fluids at rest. Therefore [we omit the arguments $(\bm{r},t)$ for simplicity],
@@
with $h=\mu+sT/n$ being the enthalpy per particle. As this relation holds for any quasiequilibrium variations of $p$, $\beta$, and $\alpha$, it can also be applied for the partial derivatives of these parameters. Since we consider a fluid at rest, the gradient of pressure vanishes due to the condition of mechanical equilibrium $\partial_{i}p=0$. Employing this, we arrive at $\partial_{i}\alpha=h\partial_{i}\beta$, and substituting this back into Eq. (110) we finally obtain
@@
Note that $U_{H}(t^{\prime})$ contains the operators of the energy density $T^{00}$ and the net particle density $N^{0}$, which do not contribute to the averages of the current operators according to Curie’s theorem. The latter states that in an isotropic medium the correlations between operators of different rank vanish de Groot et al. (1963a, b). Therefore, in the solution (A), the term $U(t^{\prime})$ can be dropped. Within the linear response approximation, we can further ignore the spacetime dependence of $\beta(\bm{r}^{\prime},t^{\prime\prime})$ in the integrand of Eq. (A) and cancel it with the term $1/\beta$.